# Weak-to-Strong Search: Align Large Language Models via Searching over Small Language Models

### v1

@@ 4.2 Chunk-level Beam Search (CBS)
The core algorithm providing the foundation of our method, Chunk-level Beam Search (CBS), is detailed in Algorithm 1 and illustrated in Figure 2. The key insight is that our beam search operates at the level of chunk. The search starts at the prompt and always maintains a hypothesis set =$\mathcal{H}=\{({\mathbf{x}},{\mathbf{y}}^{\prime})_{i}\}^{W}_{i=1}$ of $W$ states. For each state $({\mathbf{x}},{\mathbf{y}}^{\prime})$ in $\mathcal{H}$, CBS samples $K$ continuation chunks ${\mathbf{y}}_{L}$ of length $L$ from $\pi_{\text{base}}$. This results in $W\cdot K$ successor states. Among these successors, only the top-$W$ successors with the highest partial return $\log\pi^{*}({\mathbf{y}}^{\prime}\circ{\mathbf{y}}_{L}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}^{\prime}\circ{\mathbf{y}}_{L}\mid{\mathbf{x}})$ are stored in $\mathcal{H}$ and expanded further. Finally, the terminal state $({\mathbf{x}},{\mathbf{y}})$ with the highest overall return $\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$ is selected, from which the complete response ${\mathbf{y}}$ is extracted.
@@
Notably, CBS is a unified framework that encompasses several search-based algorithms: (1) CBS with $W=1$, $K=N$, $L=\infty$, is equivalent to BoN sampling with $\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$ as the scoring function. (2) CBS with $K=\infty$, $L=1$ (directly inspecting the log-likelihoods of all possible next tokens from the vocabulary) is equivalent to vanilla token-level beam search.
However, we always ensure finite chunk length and limited successor exploration via sampling (even when $\bm{L=1}$) to achieve the best of both worlds: (1) Using a finite chunk length allows CBS to discard bad states earlier and focus computational resources on expanding promising states, enhancing steerability more efficiently compared to BoN. (2) Sampling from $\pi_{\text{base}}$ with limited successor exploration implicitly enforces the KL-constraint from $\pi_{\text{base}}$ (Eq. 5b). Without this limit, integrating the KL-constraint into the objective (Eq. 5a) as in Eq. 2 would be necessary, but this can be challenging, especially when vocabularies of $\pi_{\text{base}}$ and $(\pi^{*},\pi_{\text{ref}})$ differ or with black-box language models $\pi_{\text{base}}$ whose log-likelihoods are inaccessible. In addition, it can be beneficial to keep more than one promising state $\bm{W>1}$ because the partial return only approximately correlates with the overall return (Section 4.1). While the partial return provides useful long-term guidance, we can still not strictly guarantee achieving the global optima by following the best local estimates.
Discussions on computation costs and steerability-KL tradeoff. In practice, CBS samples $W\cdot K$ continuation chunks in parallel and evaluates new states by calling $(\pi^{*},\pi_{\text{base}})$ every $L$ tokens. Larger $W\cdot K$ and smaller $L$ enhance steerability at the cost of increased computations, whereas smaller $W\cdot K$ and larger $L$ sacrifice steerability for more efficient computations. Note that high steerability, while beneficial, is not always ideal as it may lead to large KL deviation and over-optimization [16].
@@ 5 Experiments
In addition to weak-to-strong search, we evaluate several existing test-time approaches that steer large language models $\pi_{\text{base}}$ using small tuned and untuned language models $(\pi^{*},\pi_{\text{ref}})$: (1) Base: we explore regular decoding from the frozen large language model with n-shot prompting (see Appendix C.1.6 for prompt details). (2) Best-of-N Sampling (BoN)  [16, 17]: BoN uses $r=\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$ to select the highest-scoring responses among the $N$ independent responses from the frozen large language model. Since weak-to-strong search (CBS) samples $W\cdot K$ response chunks in parallel, for fair computational comparisons, we always ensure $N=W\cdot K$. (3) Emulated Fine-Tuning (EFT) [34, 12, 11, 35]: EFT approximates the results of directly fine-tuning the large language model by sampling from $\log\pi_{\text{EFT}}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})\propto\log\pi_{\text{base}}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})+\beta^{-1}(\log\pi^{*}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})-\log\pi_{\text{ref}}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t}))$, where $\beta$ is the hyperparameter from Eq. 2. Note that EFT is only applicable when all models share the same vocabulary (which is necessary for composing output distributions from different models). Whenever possible, we also compare test-time methods against directly fine-tuning the large models in the same way small models are tuned.
@@ 5.1 Controlled-Sentiment Generation & Summarization
For these two tasks, we follow the synthetic setups from [16, 46, 4], assuming access to a gold reward model $r_{\text{gold}}$. For controlled-sentiment generation, $r_{\text{gold}}$ encourages positive continuations of movie reviews, while for summarization, it encourages high-quality summaries of Reddit posts (details in Appendix C.1.4). We generate synthetic preference datasets $\mathcal{D}=\{({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})_{i}\}_{i=1}^{N}$ from $r_{\text{gold}}$ with $p({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}\mid{\mathbf{x}})=\sigma(r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{1})-r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{2}))$ to mimic human feedback [47].
@@ Results.
Figure 3 demonstrates weak-to-strong search’s great flexibility and steerability in both tasks. For summarization, weak-to-strong search consistently outperforms other test-time methods by large margins. For controlled-sentiment generation, weak-to-strong search is second only to EFT with a carefully selected hyperparameter ($\beta^{*}=1/4$) when EFT is applicable. We hypothesize that token-level adjustments from EFT are sufficient for controlled-sentiment generation, which primarily requires minor stylistic changes at the token level (e.g., “hate” $\rightarrow$ “love”). However, in the more complex task of summarization, where broader sequence-level manipulations are essential, weak-to-strong search excels. Please refer to Appendix E for quantitative comparisons of samples from different methods. We need to mention that we do not meaningfully tune weak-to-strong search (CBS)’s hyperparameters to obtain the results in Figure 3 (we use a fixed set of hyperparameters of $(4,4,5)$ for $W,K,L$ across all models), which may underestimate the performance of our method. In addition, our method enables consistent weak-to-strong generalization in the harder task of summarization: most large pre-trained models (except for gpt2-large) are stronger than the tuned gpt2 in summarizing long text, but the weak models are still able to improve the strong models through test-time guidance, nearly matching the results of direct fine-tuning. The phenomenon of weak-to-strong generalization will be further studied in Section 5.2.
Figure 3: The gold reward achieved for different large pre-trained models under the gpt2 guidance. We show the mean reward (±$\pm 2$ standard deviations) across three random seeds. EFT ($\beta^{*}$) denotes the best EFT results among $\beta\in\{1/4,1/2,1,2,4\}$; Weak-to-strong search $(4,4,5)$ denotes CBS with $W,K,L=4,4,5$; BoN ($16$) denotes BoN with $N=16$.
@@
We perform ablations to understand how CBS hyperparameters (beam width $W$, successors per state $K$, and chunk length $L$) influence performance. Figure 4 displays the ablation results for $W,K$. With the same computation budget (i.e., $W\cdot K$), the optimal trade-off between $W$ and $K$ varies by tasks: for controlled-sentiment generation, the best results come from retaining the most promising state and concentrating computational efforts on expanding from it $(W,K=1,16)$; in contrast, for summarization, maintaining multiple hypotheses $(W,K=8,2)$ yields the best results by avoiding local optima. Figure 5 displays the ablation results for $L$ where smaller $L$ benefits controlled-sentiment generation, while an intermediate $L$ is optimal for summarization. These results are consistent with our findings in Figure 3, suggesting that the simple nature of controlled-sentiment generation makes token-level manipulation sufficient and partial return a more reliable indicator of overall return. See Appendix D.1 for extended ablations.
@@
Figure 4: W, K ablations for CBS ($\mathbf{L=5}$). We show the mean rewards across three random seeds. With the same computation budget (i.e., same $W\cdot K$) the optimal hyperparameters differ by tasks.
@@
Figure 5: L ablations for CBS (W, K = 4, 4). We show the mean rewards ($\pm$ 2 standard deviations).
Figure 6: The length-controlled win rates against gpt-4-turbo for various instruction-tuned models under the Zephyr (left) or Tulu (right) guidance. Hyperparameters are in Appendix C.2.2.
@@ 5.2 Instruction Following
For small language models, we reuse two high-ranking 7B model pairs from the AlpacaEval 2.0 leaderboard as guidance: (1) Zephyr guidance: zephyr-7b-beta and its untuned version mistral-7b-sft-beta; (2) Tulu guidance: tulu-2-dpo-7b and its untuned version tulu-2-7b. All four models use the Llama-2 tokenizer. The large instruction-tuned language models we aim to further align fall into three categories: (1) same vocabulary: Llama-2-7b-chat, Llama-2-70b-chat; (2) cross vocabulary: Llama-3-8B-Instruct, Llama-3-70B-Instruct; (3) black box: gpt-3.5-turbo-instruct. As it is nearly impossible to reproduce the exact training pipeline for these small models ($\pi_{\text{ref}}\rightarrow\pi^{*}$), we do not test the baseline results of directly fine-tuning the large models as in Figure 3. Language model responses are evaluated by their length-controlled win rates (LC WR) against gpt-4-turbo, with gpt-4-turbo serving as the judge.
@@ Results.
Experimental results with Zephyr and Tulu guidance are shown in Figure 6 (detailed hyperparameters in Appendix C.2.2). Weak-to-strong search consistently outperforms other test-time baselines with great margins. There are two crucial takeaways worth mentioning: (1) Weak-to-strong search makes strong models stronger with only weak test-time guidance. Take Zephyr guidance for an example (Figure 6, left), even if most large instruction-tuned models $\pi_{\text{base}}$ are stronger than zephyr-7b-beta before steering, weak-to-strong search is still able to enhance their performances using weak models as guidance. Conversely, EFT and BoN mainly interpolate between weak and strong models, resulting in limited, if any, improvements over the strong models. We also tested beam search without external guidance [20] but we found no obvious improvements (Table 2), probably because the latent reward functions behind these language models are not well aligned with the human preference that gpt-4-turbo approximates. The same observations apply to Tulu guidance, even though the tuned tulu-2-dpo-7b is weaker than all the large instruction-tuned language models by significant margins (Figure 6, right). (2) Weak-to-strong search applies to black-box language models. Our method, requiring only sampling from large language models, is also effective for black-box models like gpt-3.5-turbo-instruct. For weak-to-strong search with gpt-3.5-turbo-instruct, we use a chunk length of 100, as the black-box APIs are stateless and do not retain activation caches, making repeated context embedding costly. Despite the long chunk length, our method effectively aligns black-box models, significantly outperforming BoN, a special case of weak-to-strong search (CBS) with infinite chunk length.
@@ 6 Discussion
We have presented weak-to-strong search, an alignment method that keeps the large language model frozen while steering its decoding through a test-time greedy search over small language models. This method builds on the insight that the log-likelihood difference between small tuned and untuned language models can serve both as a dense reward function and a critic, and introduces a novel beam search algorithm designed for balancing reward maximization and KL minimization. This method offers a compute-efficient model up-scaling strategy that eliminates the complexity of directly fine-tuning the large models, and exemplifies weak-to-strong generalization [21] that makes strong models stronger with only weak test-time guidance. Empirically, this approach is effective in controlled-sentiment generation, summarization, and instruction following.
@@
While our work focuses on aligning with human preferences, weak-to-strong search could also apply to tasks like reasoning [49, 50] and coding [51], where ground truth answers exist. This is because any pair of tuned and untuned language models can act as test-time steering forces, without necessarily being trained on preferences. This then raises several questions beyond the scope of our current study: (1) In our study, we consistently use SFTed policy as the untuned model $\pi_{\text{ref}}$ due to the two-stage nature of preference learning. In general single-stage fine-tuning tasks, does weak-to-strong search still work with a pre-trained model serving as the untuned model $\pi_{\text{ref}}$? (2) Can weak-to-strong search enhance language models in tasks where ground truth answers exist, beyond merely tailoring their knowledge and skills to human preferences?

### v3

@@ 4.2 Chunk-level Beam Search (CBS)
The core algorithm providing the foundation of our method, Chunk-level Beam Search (CBS), is detailed in Algorithm 1 and illustrated in Figure 2. The key insight is that our beam search operates at the level of chunk. The search starts at the prompt and always maintains a hypothesis set =$\mathcal{H}=\{({\mathbf{x}},{\mathbf{y}}^{\prime})_{i}\}^{W}_{i=1}$ of $W$ states. For each state $({\mathbf{x}},{\mathbf{y}}^{\prime})$ in $\mathcal{H}$, CBS samples $K$ continuation chunks ${\mathbf{y}}_{L}$ of $L$ tokens from $\pi_{\text{base}}$. This results in $WK$ successor states. Among these successors, only the top-$W$ successors with the highest partial return $\log\pi^{*}({\mathbf{y}}^{\prime}\circ{\mathbf{y}}_{L}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}^{\prime}\circ{\mathbf{y}}_{L}\mid{\mathbf{x}})$ are stored in $\mathcal{H}$ and expanded further. Finally, the terminal state $({\mathbf{x}},{\mathbf{y}})$ with the highest intermediate return $\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$ is selected, from which the complete response ${\mathbf{y}}$ is extracted. Notably, if the model to steer $\pi_{\text{base}}$ has a different vocabulary from the tuned and unturned models $(\pi^{*},\pi_{\text{ref}})$, we should first decode the sampled tokens into natural language using $\pi_{\text{base}}$’s vocabulary and then re-encode using ($\pi^{*},\pi_{\text{ref}}$)’s vocabulary for evaluation.
@@
CBS is a unified framework that encompasses several search-based algorithms: (1) CBS with $W=1$, $K=N$, $L=\infty$ (i.e., infinite chunk length) is equivalent to BoN sampling with $\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$ as the scoring function, and (2) CBS with $K=\infty$, $L=1$ (i.e., exploring all possible next tokens from the vocabulary) is equivalent to vanilla token-level beam search. However, we always ensure finite chunk length and limited successor exploration via sampling to achieve the best of both worlds: (1) Using a finite chunk length allows CBS to prune bad states during generation, enhancing steerability more efficiently compared to BoN. (2) Sampling from $\pi_{\text{base}}$ with limited successor exploration implicitly enforces the KL-constraint from $\pi_{\text{base}}$ (Eq. 5b); otherwise, integrating the KL-constraint into the objective (Eq. 5a) would be necessary for token-level search, but this can be challenging, especially when vocabularies of models differ or with black-box language base models $\pi_{\text{base}}$ whose log-probabilities are inaccessible.
Computation costs. In practice, CBS samples $WK$ continuation chunks in parallel from the frozen base model $\pi_{\text{base}}$ and prune states by calling tuned and untuned model pair $(\pi^{*},\pi_{\text{ref}})$ every $L$ tokens. Larger $WK$ and smaller $L$ enhance steerability at the cost of increased computations. Note that high steerability, while beneficial, is not always ideal as it may lead to large KL deviation and over-optimization [16].
@@ 5 Experiments
In addition to weak-to-strong search, we evaluate several existing test-time approaches that steer a large language model $\pi_{\text{base}}$ using small tuned and untuned language models $(\pi^{*},\pi_{\text{ref}})$: (1) Base: we explore regular decoding from the frozen large language model with n-shot prompting (see Appendix B.1.6 for prompt details). (2) Best-of-N Sampling (BoN)  [16, 17]: BoN uses $r=\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$ to select the highest-scoring responses among the $N$ independent responses from the frozen large language model. Since weak-to-strong search (CBS) samples $WK$ response chunks in parallel, for fair computational comparisons, we always ensure $N=WK$. (3) Emulated Fine-Tuning (EFT) [34, 12, 11, 35]: EFT approximates the results of directly fine-tuning the large language model by sampling from $\log\pi_{\text{EFT}}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})\propto\log\pi_{\text{base}}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})+\beta^{-1}(\log\pi^{*}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})-\log\pi_{\text{ref}}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t}))$, where $\beta$ is the hyperparameter from Eq. 2. Note that EFT is only applicable when all models share the same vocabulary (which is necessary for composing output distributions from different models). Whenever possible, we also compare test-time methods against directly fine-tuning the large models in the same way small models are tuned.
@@ 5.1 Controlled-Sentiment Generation & Summarization
For these two tasks, we follow the synthetic setups from [16, 46, 4], assuming access to a gold reward model $r_{\text{gold}}$. For controlled-sentiment generation, $r_{\text{gold}}$ encourages positive continuations of movie reviews, while for summarization, it encourages high-quality summaries of Reddit posts (details in Appendix B.1.4). We generate synthetic preference datasets $\mathcal{D}=\{({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})_{i}\}_{i=1}^{N}$ from $r_{\text{gold}}$ with $p({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}\mid{\mathbf{x}})=\sigma(r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{1})-r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{2}))$ to mimic human feedback [47].
@@ Results.
Figure 3 demonstrates weak-to-strong search’s great flexibility and steerability in both tasks. For summarization, weak-to-strong search consistently outperforms other test-time methods by large margins. For controlled-sentiment generation, weak-to-strong search is second only to EFT with a carefully selected hyperparameter ($\beta^{*}=1/4$) when EFT is applicable. We hypothesize that token-level adjustments from EFT are sufficient for controlled-sentiment generation, which primarily requires minor stylistic changes at the token level (e.g., “hate” $\rightarrow$ “love”). However, in the more complex task of summarization, where broader subsequence-level manipulations are essential, weak-to-strong search excels. Please refer to Appendix D for quantitative comparisons of samples from different methods. We need to mention that we do not meaningfully tune weak-to-strong search (CBS)’s hyperparameters to obtain the results in Figure 3 (we use a fixed set of hyperparameters of $(4,4,5)$ for $W,K,L$ across all models), which may underestimate the performance of our method. In addition, our method enables consistent weak-to-strong generalization in the harder task of summarization: most large pre-trained models (except for gpt2-large) are stronger than the tuned gpt2 in summarizing long text, but the weak models are still able to improve the strong models through test-time guidance, nearly matching the results of direct fine-tuning. The phenomenon of weak-to-strong generalization will be further studied in Section 5.2.
Figure 3: The gold reward achieved for different large pre-trained models under the gpt2 guidance. We show the mean reward ($\pm$ standard deviations) across three random seeds. EFT ($\beta^{*}$) denotes the best EFT results among $\beta\in\{1/4,1/2,1,2,4\}$; Weak-to-strong search $(4,4,5)$ denotes CBS with $W,K,L=4,4,5$; BoN ($16$) denotes BoN with $N=16$.
@@
We perform additional ablations to understand how CBS hyperparameters (beam width $W$, successors per state $K$, and chunk length $L$) influence performance. Figure 4 displays the ablation results for $W$ and $K$. With the same computation budget (i.e., $WK$), the optimal trade-off between $W$ and $K$ varies by tasks: for controlled-sentiment generation, the best results come from retaining the most promising state and concentrating computational efforts on expanding from it $(W,K=1,16)$; in contrast, for summarization, maintaining multiple hypotheses $(W,K=8,2)$ yields the best results probably because it helps avoid local optima. Figure 5 displays the ablation results for $L$ where smaller $L$ benefits controlled-sentiment generation, while an intermediate $L$ is optimal for summarization. These results are consistent with our findings from Figure 3, suggesting that the simple nature of controlled-sentiment generation makes token-level manipulation sufficient and cumulative reward mid-generation a more reliable indicator of overall return. See Appendix C.1 for extended ablations on more models.
@@
Figure 4: W, K ablations for CBS ($\mathbf{L=5}$). We show the mean rewards across three random seeds. With the same computation budget (i.e., same $WK$), the optimal hyperparameters differ by tasks.
@@
Figure 5: L ablations for CBS (W, K = 4, 4). We show the mean rewards ($\pm$ standard deviations) across three random seeds.
Figure 6: The length-controlled win rates against gpt-4-turbo for various instruction-tuned models under Zephyr (left) or Tulu (right) guidance. Hyperparameters are in Appendix B.2.2.
@@ 5.2 Instruction Following
For small language models, we reuse two high-ranking 7B model pairs from the AlpacaEval 2.0 leaderboard as guidance: (1) Zephyr guidance: zephyr-7b-beta and its untuned version mistral-7b-sft-beta; (2) Tulu guidance: tulu-2-dpo-7b and its untuned version tulu-2-7b. All four models use the Llama-2 tokenizer. The large instruction-tuned language models we aim to further align fall into three categories: (1) same vocabulary: Llama-2-7b-chat, Llama-2-70b-chat; (2) cross vocabulary: Llama-3-8B-Instruct, Llama-3-70B-Instruct; and (3) black box: gpt-3.5-turbo-instruct. As it is nearly impossible to reproduce the exact training pipeline for these small models ($\pi_{\text{ref}}\rightarrow\pi^{*}$), we do not test the baseline results of directly fine-tuning the large models as in Figure 3. Language model responses are evaluated by their length-controlled win rates (LC WR) against gpt-4-turbo, with gpt-4-turbo serving as the judge.
@@ Results.
Experimental results with Zephyr and Tulu guidance are shown in Figure 6 (detailed hyperparameters in Appendix B.2.2). Weak-to-strong search consistently outperforms other test-time baselines with great margins. There are two crucial takeaways worth mentioning: (1) Weak-to-strong search makes strong models stronger with only weak test-time guidance. Take Zephyr guidance for an example (Figure 6, left), even if most large instruction-tuned models $\pi_{\text{base}}$ are stronger than zephyr-7b-beta before steering, weak-to-strong search is still able to enhance their performances using weak models as guidance. Conversely, EFT and BoN mainly interpolate between weak and strong models, resulting in limited, if any, improvements over the strong models. We also tested beam search over the strong models without external guidance [20] but we found no obvious improvements compared with regular decoding (Table 2), probably because the latent reward functions behind these language models are not well aligned with the human preference that gpt-4-turbo approximates. The same observations apply to Tulu guidance, even though the tuned tulu-2-dpo-7b is weaker than all the large instruction-tuned language models by significant margins (Figure 6, right). (2) Weak-to-strong search applies to black-box language models. Our method, requiring only sampling from large language models, is also effective for black-box models like gpt-3.5-turbo-instruct. For weak-to-strong search with gpt-3.5-turbo-instruct, we use a relatively long chunk length of 100, as the black-box language model APIs are stateless and do not retain activation caches, making repeated context embedding costly. Despite the long chunk length, our method still effectively improves the alignment of black-box models, significantly outperforming BoN, a special case of weak-to-strong search (CBS) with infinite chunk length.
@@ 6 Discussion
We have presented weak-to-strong search, an alignment method that keeps the large language model frozen while steering its decoding through a test-time greedy search over small language models. This method builds on the insight that the log-probability difference between small tuned and untuned language models can serve both as a dense reward function and a value function, and then introduces a novel beam search algorithm designed for balancing reward maximization and KL minimization. This method offers a compute-efficient model up-scaling strategy that eliminates the complexity of directly fine-tuning the large models, and exemplifies weak-to-strong generalization [21] that makes strong models stronger with only weak test-time guidance. Empirically, this approach is effective in controlled-sentiment generation, summarization, and instruction following.
@@
While our work focuses on aligning with human preferences, weak-to-strong search could also apply to tasks like reasoning [49, 50] and coding [51], where ground truth answers exist. This is because any pair of tuned and untuned language models can act as test-time steering forces, without necessarily being trained on preferences. This then raises several questions beyond the scope of our current study: (1) In our study, we consistently use SFTed policy as the untuned model $\pi_{\text{ref}}$ due to the two-stage nature of preference learning; however, in single-stage fine-tuning tasks, does weak-to-strong search still work with a pre-trained model serving as the untuned model $\pi_{\text{ref}}$? (2) Although our method shows consistent weak-to-strong generalization across diverse alignment tasks, it is also critical to probe its potential failure modes [52]. Can weak-to-strong search enhance language models in tasks where ground truth answers exist, beyond merely tailoring their knowledge and skills to human preferences? (3) Additionally, while our work mainly focuses on how dense language model reward function (Eq. 4) benefits language model decoding at test time, it’s also worth exploring the potential benefits of this reward parametrization for RL tuning. Although Appendix C.3 presents some promising preliminary results, we leave further analysis for future work.
Acknowledgements
This work was supported in part by the National Key R&D Program of China (NO.2022ZD0160201). We would like to thank anonymous reviewers for their valuable feedback and helpful discussions.
Author Contributions
Zhanhui Zhou led the project, proposed the research idea, wrote the codebase, designed the experiments, conducted most of the initial experiments, and wrote the paper. Zhixuan Liu assisted with running many of the ablation studies throughout the experiments presented in the paper. All other authors provided feedback throughout the project.